OpenStudy (anonymous):

sin^2x=.5sinx sinx=.5 x=pi/6 interval: 0

OpenStudy (anonymous):

it's quadratic and should have 2 roots. tosolve you must subtract ,5sinx from both sides, factor and find the zeros.

OpenStudy (anonymous):

there are 2 roots in the interval

OpenStudy (mathmale):

sin^2x=.5sinx might be easier to understand and solve if written as 2(sin x)^2 = sin x. Solve for sin x first. And yes, there are TWO primary roots on the given interval.

OpenStudy (anonymous):

\[\sin ^2x-.5\sin x=0,\sin x \left( \sin x-.5 \right)=0\] \[Either \sin x=0,x=?\] \[or \sin x=0.5,x=?\]

OpenStudy (anonymous):

so is the error that it needs two primary roots on the interval?

OpenStudy (mathmale):

@15HReeves : In your shoes, I'd solve this problem myself. Then, I'd go back and compare my answers to the ones given. Then I'd likely quickly spot any discrepancies.

OpenStudy (mathmale):

Worry about the error later. Find the correct solutions now, and then compare your solutions with that / those given.

OpenStudy (anonymous):

\[x^2=.5x\] if you divide by \(x\) and write \[x=.5\] then you miss the obvious solution \(x=0\)